The table below reports a simple example of the U.S. balance of payments. To keep things simple, assume that the balance of payments includes only the value of goods and services, and the value of real and financial assets. Note that the table is not complete. You wiill need to calculate the balances in each account. They are not zero. Current Account (billions of dollars) Capital account (billions of dollars) Goods and Services Real and Financial Assets U.S. assets owned abroad ExportS 8000 7000 Foreign-owned assets in the United States Imports 9000 Balance Balance Assume that there are no other items in either the current account or the capital account. Given the information in the table, the value of imports to the U.S. is $ billion. The table below reports a simple example of the U.S. balance of payments. To keep things simple, assume that the balance of payments includes only the value of goods and services, and the value of real and financial assets. Note that the table is not complete. You wiill need to calculate the balances in each account. They are not zero. Current Account (billions of dollars) Capital Account (billions of dollars) Goods and Services Real and Financial Assets U.S. assets owned abroad Exports Imports Balance 3000 -1000 Foreign-owned assets in the United States 3400 Balance Assume that there are no other items in either the current account or the capital account. Given the information in the table, the value of foreign-owned assets in the United States is S billion.
Answer #1

For Balance of payment, we know that,

Current account balance + Capital account balance= 0

Capital account balance = -7000 + 9000 = 2000

So, Current account balance + 2000 = 0

or, Current account balance = - 2000

Export + Import = -2000 and Export = 8000

so, Import + 8000 = -2000

or, Import = -10000

So, given the information in the table, value of imports to the U.S. is $-10000 billion.

In the second table, in the same manner, we can find current account balance

Current Account balance = Export + Import = 3000 - 3400 = -400

Capital Account balance - 400 = 0

or, Capital Account balance = 400

or, US assets Owned Abroad + Foreign owned assets in the US = 400

or, -1000 + Foreign owned assets in the US = 400

or, Foreign owned assets in the US = 1400

So, given the information in the table, the value of Foreign owned assets in the US is $1400 billion.